the lateral area of a pyramid with a square base is 240 sq. ft. its base edges are 12 ft. long. what is the slant height of the pyramid?
Answer by ewatrrr(24785)


Hi
the lateral area of a pyramid = area of base + area of the four sides
Question states***square base, base edges =12 ft.
(240ft^2 - 144ft^2)/4 = (1/2)*lh*12
24 = (1/2)*h(L)*12
4ft = h(L)